OpenStudy (anonymous):

Can you plea solve this? (120 -6 the second power )/4+8? I keep getting a high number and my teacher's answer is 29

OpenStudy (anonymous):

you just need to make sure your paranthesis are in the right place I think. TYpe it in your calc liek this.... ((120-6^2)/4) + 8

OpenStudy (anonymous):

you will get 29 if you divide the first chunk by 4 and then add 8 to that total answer

hero (hero):

\[\frac{120-6^2}{4} + 8 = \frac{120 - 36}{4} + 8 = \frac{84}{4} + 8 = 21 + 8 = 29\]